# Out-of-Distribution Runtime Adaptation with Conformalized Neural Network Ensembles

### v1

@@ X-F Statistical Analysis for Data Collection
As mentioned in Section VII, since each of the trajectories are sampled i.i.d. from the relevant probability distribution, $\mathcal{D}$, the distribution of coverage for a threshold calibrated with conformal prediction follows a Beta distribution; this can be used to inform the data-gathering process when collecting data for a new application. Although the coverage value for an outlier detection threshold calibrated using our conformal method is the same in expectation for any valid calibration set, the empirical coverage value attained by an instance of the calibration method is subject to a probability distribution with a shape that depends on the size of the calibration set.
@@
With the value of $K$ set by $N$ and the desired coverage rate such that $K:=\lceil(1-\delta)(N+1)\rceil$, the probability of $\delta_{e}$ resolving close to its ideal value, $\delta$, increases as the value of $N$ increases. Given that the empirical coverage is Beta distributed, as described in Theorem 1 and seen in Figure 11, we can use beta functions and the corresponding shape parameters $(a,b)=(K,N+1-K)$ to calculate the probability of $\delta_{e}$ landing within a desired range of values. Specifically, we can use the Beta function, $B(a,b)$, and Incomplete Beta function evaluated at $x$, $I_{x}(a,b)$, to calculate the probability of $(1-\delta_{e})$ resolving between $x_{1}$ and $x_{2}$ for a given $N$ and $K$:
@@
(Note that this is equivalent to subtracting two instances of the cumulative distribution function for the Beta distribution.) For example, for the values $N=1000$ and $K=961$ (corresponding to $(1-\delta)\approx 0.96$), the probability of $(1-\delta_{e})$ resolving between $0.95$ and $0.97$ after a single instance of the training and calibration process is of approximately $89.65\%$. The procedure for calculating this probability for a calibration dataset containing $N_{test}$ data points is described in Algorithm 2.
@@
This principle can also be used before any data is collected, to estimate the number of calibration data points required to ensure that the empirical coverage value falls within specified bounds around the desired coverage, with a user-specified confidence (i.e. the size of the dataset required for $(1-\delta_{e})$ to be within a certain distance of $(1-\delta)$, with a desired probability, after a single iteration of the calibration procedure). A corresponding value of $N$ can be identified using a bisection search method specially adapted for this problem, described in Algorithm 3.

### v3

@@ C.12 Statistical Analysis for Data Collection
As mentioned in Section 7, since each of the trajectories are sampled i.i.d. from the relevant probability distribution, $\mathcal{D}$, the distribution of coverage for a threshold calibrated with conformal prediction follows a Beta distribution; this can be used to inform the data-gathering process when collecting data for a new application. Although the coverage value for an outlier detection threshold calibrated using our conformal method is the same as expected for any valid calibration set, the empirical coverage value attained by an instance of the calibration method is subject to a probability distribution with a shape that depends on the size of the calibration set.
@@
With the value of $K$ set by $N$ and the desired coverage rate such that $K:=\lceil(1-\delta)(N+1)\rceil$, the probability of $\delta_{e}$ resolving close to its ideal value, $\delta$, increases as the value of $N$ increases. Given that the empirical coverage follows a Beta distribution, as described in Theorem 5.2 and seen in Figure 9, we can use Beta functions and the corresponding shape parameters $(a,b)=(K,N+1-K)$ to calculate the probability of $\delta_{e}$ landing within a desired range of values. Specifically, we can use the Beta function, $B(a,b)$, and Incomplete Beta function evaluated at $x$, $I_{x}(a,b)$, to calculate the probability of $(1-\delta_{e})$ resolving between $x_{1}$ and $x_{2}$ for a given $N$ and $K$:
@@
(Note that this is equivalent to subtracting two instances of the cumulative distribution function for the Beta distribution.) For example, for the values $N=1000$ and $K=961$ (corresponding to $(1-\delta)\approx 0.96$), the probability of $(1-\delta_{e})$ resolving between $0.95$ and $0.97$ after a single instance of the training and calibration process is approximately $89.65\%$. The procedure for calculating this probability for a calibration dataset containing $N_{test}$ data points is described in Algorithm 2.
@@
This principle can also be used before any data is collected, to estimate the number of calibration data points required to ensure that the empirical coverage value falls within specified bounds around the desired coverage, with a user-specified confidence (i.e., the size of the dataset required for $(1-\delta_{e})$ to be within a certain distance of $(1-\delta)$, with a desired probability, after a single iteration of the calibration procedure). A corresponding value of $N$ can be identified using a bisection search method specially adapted for this problem, described in Algorithm 3.